2y^2+5xy=1...............equation (1)
3x-2y,=-6.....................equation (2)

Answer by Boreal(15235)

2y^2+5xy=1
3x-2y=-6; 3x=2y-6;x=(2y/3)-2
2y^2+{5(2y/3)-10}y=1; 2y^2+(10/3)y^2-10y-1=0
(16/3)y^2-10y-1=0
16y^2-30y-3=0
y=(1/32)(30+/- sqrt (900+192))
y=(1/32)(30+/-33.05)
y=(1/32)63.05)=1.970
y=(1/32)(-3.05)=-0.0953
put into second equation
3x-3.94=-6
3x=-2.06
x=-0.687
try both in top equation
7.76+5(-.687*1.97)=7.76-6.767=0.99 within range with rounding error.
(-0.687,1.97)
try second
3x+0.1906=-6
3x=-6.1906
x=-2.064
check in top equation
0.0181+(5*-2.064*-0.0953)=0.0181+0.9832 (rounding only here)=1.0013
(-2.064,-0.0953)